OpenStudy (allieatthedisco):

convert 37.5 g/mL to kg/L

hero (hero):

Hint: m = .001, k = 1000

hero (hero):

I'll be back in a sec

hero (hero):

Okay, basically change the milli and the kilo to their actual numerical values first

hero (hero):

\[\frac{37.5 g}{.001 L} = \frac{\text{____} 1000 g}{L}\]

hero (hero):

Actually, you can probably make that blank an x

hero (hero):

After cross multiplication and manipulation, you'll end up with: \[\frac{37.5 g}{x 1000 g} = \frac{.001 L}{L}\]

hero (hero):

Now, to cancel out those annoying variables that make things more difficult than they have to be: \[\frac{37.5}{x 1000} = \frac{.001}{1}\]

hero (hero):

Now, .001 is the same as 1/1000, so we can re-write the other side as: \[\frac{37.5}{x 1000} = \frac{1}{1000}\]

hero (hero):

Cross multiplication and manipulation leads us to this: 37.5 = x

hero (hero):

You probably did not understand anything I did

OpenStudy (lgbasallote):

dimension analysis method is simpler i think

hero (hero):

Maybe, but this works too.

hero (hero):

It depends on the person. Dimensional analysis is useless for me

hero (hero):

Thus, 37.5 g/mL = 37.5 kg/L

OpenStudy (allieatthedisco):

Yeah, I'll probably just try Dimensional analysis. But thanks to all who contributed! (:

OpenStudy (xishem):

Dimensional analysis is tons simpler and more dynamic: \[37.5\frac{g}{mL}\times \frac{1kg}{1000g} \times \frac{1000mL}{1L}=37.5\frac{kg}{L}\]

hero (hero):

I don't know about the "more dynamic" part.

hero (hero):

If you look at what I did more closely, you'll realize how simple my way is as well.

hero (hero):

The only reason my steps appear longer is because of my attempts to explain the methods I used.

OpenStudy (anonymous):

I would just like to say that I too think dimensional analysis is not a more effective way to solve these problems. I find that dimensional analysis makes thing more complicated than they should be.

OpenStudy (anonymous):

I know this does not actually answer the question, but I just felt the need to input my opinion.

OpenStudy (xishem):

How do you deal with converting complicated products of units with your system? Like:\[k_e=8.988 \times 10^9 N \cdot m^2 \cdot C^{-2}\]Besides setting up several proportions.

hero (hero):

You didn't tell me what to convert it to

hero (hero):

It's not impossible

OpenStudy (anonymous):